OpenStudy (anonymous):

Find the volume of a cylinder with a diameter of 8 inches and a height that is three times the radius. Use 3.14 for pi and round your answer to the nearest hundredth.

OpenStudy (anonymous):

sleek-feathered one

OpenStudy (jdoe0001):

volume of a cylinder = \(\large \pi r^2 h\) r = radius, h = height

OpenStudy (anonymous):

I know that. I need help with the height part. Is it 5*3 Or 5^3

OpenStudy (jdoe0001):

well the diameter is 8... and the radius is half of the diameter thus \(\bf radius=\frac{\cancel{ 8 }}{\cancel{ 2 }}\to 4\qquad {\color{brown}{ 3}}\times 4\to 12=h\)

OpenStudy (anonymous):

Opps, sorry I meant 4. But thanks!